OpenStudy (anonymous):

Find all complex solutions of 7x^2+2x+3=0

OpenStudy (anonymous):

@Hero @ParthKohli @whpalmer4

OpenStudy (whpalmer4):

do the same thing you did to solve the previous problems. the only difference is that you'll have a square root of a negative number, and \[\sqrt{-x} = i\sqrt{x}\]where \(i = \sqrt{-1}\)

OpenStudy (anonymous):

All I got was there was no solutions.

OpenStudy (whpalmer4):

how did you solve it?

OpenStudy (anonymous):

Completing the square. What I got now was x=1/7(-1+2isqrt5) and x=1/7(-1-2isqrt5)